the river current is 8 mph. a jet ski can travel 34 miles downstream in the same time that it takes to go 18 miles upstream. what is the speed of th jet ski with no current?
Answer by Alan3354(30993)

the river current is 8 mph. a jet ski can travel 34 miles downstream in the same time that it takes to go 18 miles upstream. what is the speed of th jet ski with no current?
--------------------
c = 8
j = jet ski's speed
-------------
d = rt
t = d/r
t = 34/(j+c) = 18/(j-c)
34/(j+8) = 18/(j-8)
34(j-8) = 18(j+8)
34j-272 = 18j+144
16j = 416
j = 26 mph